the balls are released from top of the tower of height H at regular interval of time. When first ball reaches the ground, the nth ball is to be just released and (n+10)/2 th ball is at some distance 'h' from the tp of the tower. Find the value of h.

Grade:11

1 Answers

Arun
3 years ago
Let the balls are release at time interval t
At time t second ball is released, at time 2t from the start, 3rd ball is released
Therefore at time , the th ball will be released
And, at time (n-1)t, the nth ball will be released
Time taken in reaching the first ball to the ground
using the second equation of motion
H = (1/2)gT²
T = (n - 1)t
Thus
or, 
We have to find the distance travelled by th ball in time
(n-1)t - (n-1)t/2= (n-1)t/2
The distance travelled = h
 
 
Therefore, the ball is at distance H/4 from the top of the tower.